trying to make sense and solve the following:
3h-x=2(l+t);solve for h

Answer by richwmiller(17219)

To solve for h we need to get h on one side all by itself.
Fortunately that will be easy
3h=2*(l+t)+x
h=(2*(l+t)+x)/3
h=1/3*(2l+2t+x)